# Quantum mechanics provides the physical basis of teleological evolutions

### v1

@@ 2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm. We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of a quantum algorithm, all the information is selected by the initial measurement. Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. That performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. As we can see in Section (3) of the Appendix, in this new initial measurement outcome it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps. This is the physical correspondent of one of the logical causal loops of the advanced knowledge rule.
@@ 2.4 Second demonstration of the advanced knowledge rule
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution. In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
@@ 2.5 Origin of the retrocausality
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^3] By the way, classical logic is neutral for what concerns the notion of causality. Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept. In principle, classical logic should also go along with the coexistence of causality and retrocausality.. This said, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
Summarizing, at the root of the retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it. By the way, note that this is not an idle speculation, as might even seem. Indeed, we are dealing with a retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
@@ 2.6 Generalization
The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup. However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables. Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says. Of course, in the case that we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule. The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops which locally explain it and play the role of the hidden variables envisaged by Einstein et al [9]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5]. Note that this is what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
@@
In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it. Of course, it also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
This ends the first part of the work. The second is about the consequences of the advanced knowledge rule or the quantum correlation time-symmetrization rule. These consequences are both widespread and destabilizing: indeed, the time-symmetric form of retrocausality implicit in these rules yields a different way of viewing physical reality.
@@
The consequence just seen, demonstrating the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [9], is destabilizing for the following reason. After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [10] and Ekert’s quantum cryptography based on it [11], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
Another destabilizing consequence of the advanced knowledge rule is that, according to it, quantum algorithms are teleological evolutions, whereas, currently, science claims that teleological evolutions are without physical basis.
An even more destabilizing consequence is that the physical basis of the teleological character of quantum algorithms, under the quantum cosmological assumption, applies as well to the evolutions of the living. Of course, the teleological notion was once applied by natural philosophy to the evolution of life; by natural philosophy, it is also applied today [12]. As already mentioned, with the advent of modern science, it was dismissed for the alleged absence of a physical basis. For its finalistic character, it has also been ascribed to the category of beliefs. The next subsection is dedicated to the teleological notion under the present theory of the quantum computational speedup.
2.7 The physical basis of teleological evolutions
Let us repeat another time the advanced knowledge rule, which naturally implies the teleological character of quantum algorithms.
We have demonstrated this rule in two redundant ways. The first, by describing the quantum speedup in a classical logic way, the second by a time-symmetrization of the ordinary quantum description of the quantum algorithm which is mandatory to satisfy the atemporal symmetry implicit in it: we have seen that this time-symmetrization generates the physical correspondent of the advanced knowledge rule. With this, of course, we have also demonstrated the teleological character of quantum algorithms: indeed, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
We have also seen that, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm. As such, it can occur with a quadratic quantum speedup. The “can occur” is in the sense that there is certainly room for a quantum speedup, but the unitary evolution must benefit of it; it should occur with a just suboptimal quantum speedup if the search for the value of life is random [21] .
We go now to the fact that this possible teleological character would be inherited by the evolution of life on Earth and the evolution of species.
@@ 2.7 The physical basis of teleological evolutions
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life applies to the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution is derivable from the former evolution by tracing out all the rest. Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former. In summary, the evolution of life on Earth could occur with quantum speedup (in the sense that there is certainly the room for that) and consequently be teleological in character; it should actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
@@
All this is obviously destabilizing, it is a way of viewing the macroscopic reality that goes against the modern way of viewing it, and revamps ancient natural philosophy views which were ascribed by science to completely unscientific beliefs.
Let us then discuss more extensively this contrast between the scientific and the natural philosophy view of the teleological notion.
The notion of teleology as final cause was introduced almost two and a half millennia ago by Plato and Aristotle to explain the “formation” of the functional structure of the living (Aristotle: teeth were created for the purpose of chewing). Of course, at the time, there was not yet the notion of evolution of the living. In more recent times, the teleological notion was applied by natural philosophy to the evolutions of the living—by natural philosophy it is also applied today [12]. It was then dismissed by science for the alleged absence of a physical basis—precisely because it implies a retrocausality deemed to be physically impossible. Of course, at the time of classical physics, this was completely justified. For its finalistic character, the teleological notion was also ascribed to the category of beliefs.
It should however be noted that this dismissal was not without regret; indeed, it meant doing without a possibly enlightening relation between the notion of teleological evolution of life and that of the functional structure of living (teeth were created for the purpose of chewing). We cite two authoritative objections to this dismissal. One is a famous sentence by the evolutionary biologist Haldane [26]:
@@
We quote a precedent in the quantum description of macroscopic natural evolutions. It is about the asymptotic behavior of predator species competing for the same prey [27]. It is explained by a quantum model, whereas it would be difficult to explain it in a classical way. In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing out all the rest. Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way. In other words, we should accept quantum superposition, entanglement, and correlation between the observables of macroscopic objects.
@@
By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter. However, naturally, classical physics turns out to work extremely well in describing macroscopic reality. Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions. This situation might point to a particular object of research. It could be interesting to study how the operation of “tracing-over the rest” (at the toy model level) reconciles a teleological behavior in the long term with a classical behavior in the short term.
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, it would be an even more fundamental discovery than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [15] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—an attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [16] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [17]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [18]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [19] in 1985.
@@ 3 Precedents
It is a work by Costa de Beauregard [22] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar. Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years. Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory. See [23] for an account of this episode. It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [24, 25]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined. On the contrary, the present time-symmetric form of retrocausality should be sort of universal for quantum correlations. The possible relation of those works to the present work, as things are now, would not be evident. Of course, it could be the object of further investigation.

### v2

@@ 2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm. We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of the quantum algorithm, all the information is selected by the initial measurement. Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, that performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum algorithm. As we can see in Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum algorithm it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps. This is the physical correspondent of one of the logical causal loops whose mutually exclusive or is logically implied by the advanced knowledge rule.
@@ 2.4 Second demonstration of the advanced knowledge rule
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution. In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
@@ 2.5 Origin of the retrocausality
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality. Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept. In principle, classical logic should also go along with the coexistence of causality and retrocausality.. Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it. By the way, note that this is not an idle speculation, as might even seem. Indeed, we are dealing with a logical form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
@@ 2.6 Generalization
The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup. However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables. Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says—see the Appendix. Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule. The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5]. Note that this is what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
@@
In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it. Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
@@
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [15] might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolution of the universe toward the value of life. We repeat here for convenience the reason for its possible quantum speedup and consequent teleological character.
Under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life, in the other universes where this value is different there is absence of life.
We have seen in the Generalization subsection that any such building up of quantum correlation can always occur with quantum speedup in the sense that there is certainly the room for it. It can consequently be teleological in character.
@@ 2.7 About the teleological character of the evolutions of the living
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life holds for the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from the former evolution by tracing all the rest. Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
In summary, under the quantum cosmological assumption, the evolution of life on Earth can occur with quantum speedup and consequently be teleological in character; it might actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
@@
All this is obviously destabilizing; it is a way of viewing the macroscopic reality that goes against the current (scientific) way of viewing it, and revamps the natural philosophy views that were ascribed by science to the unscientific beliefs. Of course, the fact that what occurs to us today is affected by the drying up of an area of Africa three millions of years ago is obvious, but that it is also affected by what we will be in three million years’ time, to the current way of thinking is more difficult to digest.
@@
We quote a precedent in the quantum description of macroscopic natural evolutions. It is about the asymptotic behavior of predator species competing for the same prey [18]. It is explained by a quantum model, whereas it would be difficult to explain it in a classical way. In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing-out all the rest. Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way. In other words, we should accept quantum superposition, entanglement, and correlation between the observables of macroscopic objects.
@@
By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter. However, naturally, classical physics turns out to work extremely well in describing macroscopic reality. Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions. This situation might point to a particular object of research. It could be interesting to study how the operation of “tracing-out all the rest” (at the toy model level) reconciles the teleological character of an evolution in the longer term with its classical character in the shorter term. This could also be a further perspective under which to address the problem of unifying classical and quantum theory.
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, this discovery would be even more fundamental than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [19] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of an attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [20] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [21]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [22]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [23] in 1985.
@@ 3 Precedents
It is a work by Costa de Beauregard [24] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar. Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years. Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory, which also implied a form of retrocausality. See [25] for an account of this episode. It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [26, 27]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined. On the contrary, the present time-symmetric form of retrocausality is sort of universal for quantum correlations. The possible relation of those works to the present work, as things are now, would not be evident. Of course, it could be the object of study.